OpenStudy (anonymous):

What is the slope between the two points (4, 14) and (2, 4)?

hero (hero):

Hint: Set up two equations using y = mx + b

OpenStudy (anonymous):

Okay, thank you so much!

hero (hero):

14 = 4m + b 4 = 2m + b Isolate b in both equations: b = 14 - 4m b = 4 - 2m

hero (hero):

set b = b and solve for m 14 - 4m = 4 - 2m 14 - 4 = 4m - 2m 10 = 2m 10/2 = m 5 = m Therefore, the slope is 5

hero (hero):

All you need is y = mx + b

hero (hero):

Remember, in the formula y = mx + b, m represents the slope.